OpenStudy (anonymous):

For a regular n-gon: a. What is the sum of the measures of its angles? b. What is the measure of each angle? c. What is the sum of the measures of its exterior angles, one at each vertex? d. What is the measure of each exterior angle? e. Find the sum of your answers to parts b and d. Explain why this sum makes sense. can you Please explain

Directrix (directrix):

Are you talking the sum of the interior or exterior angles of the regular n-gon?

OpenStudy (anonymous):

yep

Directrix (directrix):

not for this: a. What is the sum of the measures of its angles?

OpenStudy (anonymous):

(n-2)*180/n

OpenStudy (anonymous):

i gave this as an answer

OpenStudy (anonymous):

N-gon Ext angle = 360/ N-gon Int angle = 180 - 360/N-gon Ratio is 360/N-gon : 180 - 360/N-gon Ratio is 360/N-gon : (180*N-gon - 360)/N-Gon Ratio is 360 : 180*N-gon - 360 Ratio = 360 : 180(N-gon - 2) Ration = 2 : 1*(N-gon - 2) Ratio = 2: (N-gon - 2)

OpenStudy (anonymous):

measure of an angle

Directrix (directrix):

That is the sum of the interior angles of any convex polygon, regular or not.

OpenStudy (anonymous):

i just dont know how to give the answer as for each itemized : a, b, c, d, e

OpenStudy (anonymous):

to find sum it is just (n-2)*180

Directrix (directrix):

b. What is the measure of each angle? [180*(n-2) ]/n for one interior angle of a regular polygon

Directrix (directrix):

@LilySwan c. What is the sum of the measures of its exterior angles, one at each vertex? You should know this from your theorems. What is it.

OpenStudy (anonymous):

ok for a the answer will be

Directrix (directrix):

C, the answer for c

OpenStudy (anonymous):

iam confused

OpenStudy (anonymous):

can you please start explaining from the option A

Directrix (directrix):

The answer to part A is a theorem.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you please indicate towards it

OpenStudy (anonymous):

i got this wrong to

Directrix (directrix):

To find sum it is just (n-2)*180 This is from @SerikMB above in the thread.

Directrix (directrix):

That is where n is the number of sides of the polygon.

Directrix (directrix):

That, you just learn - memorize.

OpenStudy (mertsj):